Fred, Tiger and Nancy all play a round of golg togehter. their combined score was is 223. Freds score was 3 more than Nancy, and tigers score was 7 more than freds. what was each persons score?
Answer by rfer(16322)

f=n+3
n=n
t=(n+3)+7
n+3+n+n+3+7=223
3n=210
n=70
n+3=73
(n+3)+7=80